How do you evaluate -6.9-(-7.7)+25?
Answer by nyc_function(2741)

-6.9 -(-7.7)+ 25 =


-6.9 + 7.7 + 25 =


-6.9 + 32.7 = 25.8


By the way, this is not an evaluation. This problem and others like it can be

solved using the rules for PEMDAS.